two regions of the human body have a temperature at 37*c and 35*c respectively. how many times their radiant emittances differ?


1
Expert's answer
2021-05-03T10:28:05-0400

"n=\frac{L_1}{L_2}=\frac{\sigma T_1^4}{\sigma T_2^4}=\frac{T_1^4}{T_2^4}=\frac{310^4}{308^4}=1.026."